# Enhancing Few-shot Keyword Spotting Performance through Pre-Trained Self-supervised Speech Models

### v1

@@ Abstract
Keyword Spotting plays a critical role in enabling hands-free interaction for battery-powered edge devices. Few-Shot Keyword Spotting (FS-KWS) addresses the scalability and adaptability challenges of traditional systems by enabling recognition of custom keywords with only a few examples. However, existing FS-KWS systems achieve subpar accuracy at desirable false acceptance rates, particularly in resource-constrained edge environments. To address these issues, we propose a training scheme that leverages self-supervised learning models for robust feature extraction, dimensionality reduction, and knowledge distillation. The teacher model, based on Wav2Vec 2.0 is trained using Sub-center ArcFace loss, which enhances inter-class separability and intra-class compactness. To enable efficient deployment on edge devices, we introduce attention-based dimensionality reduction and train a standard lightweight ResNet15 student model. We evaluate the proposed approach on the English portion of the Multilingual Spoken Words Corpus (MSWC) and the Google Speech Commands (GSC) datasets. Notably, the proposed training method improves the 10-shot classification accuracy from 33.4% to 74.1% on 11 classes at 1% false alarm accuracy on the GSC dataset, thus making it significantly better-suited for a real use case scenario.
@@ I Introduction
Keyword Spotting (KWS), which aims to detect specific spoken commands or wake-words in audio streams, has become a cornerstone technology for a wide array of devices enabling hands-free interaction. Traditional KWS systems, while effective for fixed vocabularies, are limited in adaptability and scalability, as they need to be trained thousands of training examples per keyword and require computation and memory resources beyond the capabilities of embedded edge devices, which are often battery powered smart sensors. Few-shot keyword spotting systems (FS-KWS) address these challenges by enabling the recognition of custom keywords from only a handful of examples, using deep neural networks up to only a few hundred thousand parameters.
@@
The core idea behind FS-KWS is to leverage metric learning frameworks so that the model can readily adapt to new classes (i.e. keywords) by enrolling only a few examples [1]. Prototypical networks have become the popular approach, which involves training a representation model that processes speech segments into fixed dimensional feature vectors, which are also referred to as embeddings [2]. The goal is to learn an embedding space in which samples of the same keyword cluster together, while those of different keywords are well separated. This is achieved by metric learning objectives such as Triplet loss or Prototypical Loss [3], where the latter is based on prototypes calculated within each mini-batch. For systems that support the open-set setting, users provide a few spoken examples of their keywords during the enrollment phase, from which class prototypes are computed as the mean of the embeddings [4]. Inference is carried out by calculating the distances between the embedding of the test sample and the enrolled class prototypes; and the decision is based on the shortest distance. Unfortunately, the leading approaches [5, 6] in this new field do not provide decent user experience, achieving low accuracies at desired false acceptance rates.
@@
A common approach to boost the performance of supervised speech tasks is to integrate task-agnostic self-supervised learning (SSL) models. These SSL models output a feature vector for each frame of speech, resulting in many vectors per the duration of a keyword. To obtain lower dimensionality embeddings, [7] uses simple mean-pooling followed by a linear transformation for multiple tasks including KWS; [8] proposes a two-layer convolutional encoder for KWS; and [9] demonstrates that attention-based models are useful for other tasks such as speaker identification or emotion recognition.
@@
In this paper, we propose a FS-KWS model training scheme as shown in Fig. 1(a) that outperforms leading approaches by utilizing a pretrained SSL model in tandem with an alternative metric learning objective. Specifically, our training scheme leverages knowledge distillation, where the teacher network is based on a more capable SSL model and uses Sub-center ArcFace (SCAF) loss [10] as the objective function, as opposed to triplet loss variants. ArcFace [11] was originally introduced to improve face recognition performance by enforcing inter-class separability and intra-class compactness. SCAF enhances this by introducing multiple sub-centers per class and selecting the nearest one during training, thus accommodating intra-class variance more effectively. Although initially designed for face recognition, ArcFace has recently been applied to tasks such as emotion recognition [12] and speaker verification [13]. To the best of our knowledge, our work is the first to employ SCAF in the context of audio discrimination for FS-KWS.
We also explore training strategies for the edge-friendly student model by investigating combinations of knowledge distillation losses. Using public datasets, we provide comparisons with a baseline FS-KWS approach tailored for edge deployment.
@@
  • Utilization of SCAF to efficiently reduce the dimensionality of high-dimensional embeddings from a pre-trained SSL model.
@@ II-A The Teacher Representation Model
Figure 2: Proposed Teacher Representation Model with (a) linear projection, (b) attention encoder DR models
Figure 3: Performance of (a) Teacher models on the MSWC dataset under 1-shot setting for different DR architectures (b) Student models on the MSWC dataset under 1-shot setting (c) Student models on the GSC dataset under 10-shot setting
1) Linear Projection given in Fig. 2(a) implements a straightforward dimensionality reduction approach, where temporal averaging is followed by linear projection to obtain lower-dimensional representations.
2) Attention Encoder given in Fig. 2(b) incorporates a Scaled Dot Product Attention [16] mechanism followed by a one-dimensional convolutional layer (Conv1D). The attention mechanism computes temporal relationships across the input sequence, enabling the model to focus on salient temporal features. The subsequent Conv1D layer reduces the temporal dimension to a singular value through weighted averaging, effectively preserving critical temporal information while reducing computational complexity.
@@
These dimensionality reduction (DR) models are trained using labeled samples with two distinct loss functions: (i) Triplet and (ii) SCAF. The experiments carried out to determine the optimal DR model and the loss function to train the lightweight edge model are detailed in Section III.
@@ II-B Edge-friendly Student Representation Model
ResNet-based models are widely utilized as benchmarks in edge-focused keyword spotting (KWS) approaches [5, 17, 18, 19]. Consequently, our experiments focus on training a small-footprint model using the ResNet-15 architecture. This model processes mel-frequency cepstral coefficients (MFCC) extracted from the original audio, which compresses the input data size significantly. This reduces memory usage and computational demands, hence making it well-suited for deployment on edge hardware.
@@
In this work, we employ mean squared error (MSE) as the distillation loss to align the embeddings of the teacher and student models. For task-specific loss, we evaluate two alternatives: Triplet and SCAF losses as detailed in Section III.
@@ III-A Datasets
To train the dimensionality reduction network of the teacher, and the edge-friendly student network, we use the train split of the English portion of Multilingual Spoken Words Corpus (MSWC) [21], which contains 5.3 million one-second samples, belonging to 38,150 unique words. We test our approach using the test split of the MSWC dataset, comprised of 700k samples from 38,066 unique words, in addition to the Google Speech Commands (GSC) dataset [22], a collection of 105k one-second audio clips belonging to 35 short commands.
@@ III-B Evaluation Protocol
To evaluate the few-shot keyword spotting performance, we adopt the same testing protocol as in [5], both for teacher and student models. During the enrollment phase, $K$ samples for each keyword is fed to the frozen representation model, and the mean of the generated embeddings of each keyword class is computed to generate class prototypes. For inference, the test sample is fed to the feature extractor, and the cosine distance between the test sample embedding and the prototypes of each keyword class is computed. If the closest of these distances is less than a threshold $T$, the test sample is assigned to the class with the nearest prototype, otherwise it is assigned to the ”others” class. By changing the threshold $T$, the tradeoff between accuracy levels and false acceptance rate (FAR) of ”others” can be analyzed.
@@
For the GSC dataset, we adapt the original KWS-12 test setting [22] for FS-KWS. The goal is to decide between 11 classes (target keyword classes: {on, off, left, right, up, down, go, stop, yes, no} and silence), while ideally rejecting the 25 remaining classes as ”others”. The silence class includes recordings with different types of background noise. It should be noted that the inclusion of the silence class makes our evaluation protocol more challenging compared to the baseline work of [5]. For enrollment, $K$ samples from each keyword class in the training set are used to generate the class prototypes, and the entire test set tested against these prototypes to evaluate the performance. To reduce the variance of the results due to the selection of the few-shot samples to compute the prototype, we repeat every test 100 times and average the obtained statistics.
@@ III-C Selection of Dimensionality Reduction Model
To identify the optimal teacher model for training the lightweight representation model, we conducted experiments utilizing two dimensionality reduction (DR) architectures alongside various loss functions. In our initial experiment, we assessed the performance of the two architectures, both of which were trained using triplet loss. Each model underwent triplet training for 3000 batches, with each batch sampling 512 triplets from randomly selected word samples. The squared normalized Euclidean distance was computed for each pair of samples. We set the triplet margin to 0.5 and the model checkpoint that minimized the loss on the validation set was selected for subsequent testing. As anticipated, the attention-based DR architecture yielded superior results. Consequently, we focused our investigation on the selection of loss functions specifically for this architecture. We repeated the experiment, substituting the loss function with Subcenter ArcFace loss, which employed an angular margin of m=28.6, a scale factor of s=32, and utilized K=3 subcenters per class, training for a total of 10 epochs.
@@
Fig. 3(a) illustrates the performance of three different models on the MSWC dataset under 1-shot setting. The results indicate that the attention-based dimensionality reduction architecture significantly outperforms others when trained with SCAF loss. For completeness, we trained student models using all three models as teachers. The attention-based architecture with SCAF loss also led to the best results for the student models. For the scope of this letter, we only provide results on the student models trained using this teacher model in the next section.
@@
TABLE I: Comparison of teacher and student (edge) network performance across GSC and MSWC datasets for {1, 5, 10}-shot settings.
Model Objective # of supports GSC MSWC
$ACC^{+}_{1\%}$ $ACC^{+}_{5\%}$ AUC $DET_{1\%}$ $DET_{5\%}$ AUROC
Teacher SCAF 1-shot 69.0 77.1 83.4 91.6 97.4 99.3
Edge (Resnet15) Triplet [5] 22.4 53.5 70.7 71.7 88.5 97.5
SCAF 12.4 36.1 55.9 76.1 88.1 92.1
KD 42.1 59.8 73.4 83.5 93.0 98.1
KD + Triplet 27.8 42.3 62.0 84.5 95.4 98.9
KD + SCAF 33.4 49.1 67.6 76.5 90.7 96.3
Teacher SCAF 5-shot 81.6 84.8 89.9 96.8 99.2 99.8
Edge (Resnet15) Triplet [5] 32.2 70.2 84.4 89.3 97.0 99.3
SCAF 40.5 59.0 74.2 89.1 94.6 96.5
KD 68.7 80.2 88.1 94.7 97.9 98.7
KD + Triplet 48.4 64.0 79.8 94.1 98.7 99.6
KD + SCAF 56.5 69.8 83.4 89.1 96.3 99.4
Teacher SCAF 10-shot 82.2 85.2 90.9 96.9 99.3 99.9
Edge (Resnet15) Triplet [5] 33.4 76.2 87.5 91.4 97.5 99.4
SCAF 47.5 65.1 77.8 91.0 96.1 97.4
KD 74.1 83.7 90.1 95.5 98.3 99.6
KD + Triplet 56.5 70.9 83.8 94.8 99.0 99.7
KD + SCAF 63.7 73.3 86.1 91.6 97.2 99.1
III-D Comparative Analysis Edge Model Training Strategies
@@
To systematically evaluate the training strategies, we trained the edge-friendly ResNet15 model using four approaches and compared these to the baseline method [5], wherein the same ResNet15 was trained with Triplet loss (Triplet). We directly used the trained checkpoint from the authors’ repository [^1] https://github.com/mrusci/ondevice-learning-kws to evaluate the baseline method with our testing protocol. For the input to the model, we use the first 10 MFCCs, computed using a frame size of 40 ms, a stride of 20 ms, and applying Hamming windowing. This results in a feature map of size 49x10 for each 1-second test sample.
@@
The first strategy involves the adoption of SCAF loss in place of Triplet loss, as defined in the baseline method (SCAF). The second strategy employs knowledge distillation from the selected teacher model discussed in the previous section, with the following configurations: (1) KD: trained solely with $L_{KD}$ ($\lambda=0$); (2) KD + Triplet: trained with both $L_{KD}$ and Triplet loss ($\lambda=0.03$), thereby balancing MSE and Triplet losses; (3) KD + SCAF: trained with $L_{KD}$ and SCAF loss ($\lambda=0.003$).
@@
Fig. 3(c) presents accuracy versus FAR for the GSC dataset for the 10-shot setting. The KD model consistently outperforms all other configurations at FAR = 1% and 5%. The baseline Triplet model ranks second at FAR = 5% but experiences a significant decline at lower FARs. This is primarily due to the inclusion of the silence class, as we observed in keyword-specific analysis not shown in this letter. We hypothesize that our approach, which includes knowledge distillation from the pre-trained SSL model, is critical in providing robustness for unseen conditions.
A comprehensive evaluation of the training strategies is shown in Table I, which includes the best performing teacher model and all edge-friendly model training alternatives. On the MSWC dataset, KD + Triplet strategy shows marginally better performance than KD with higher AUC and detection rates at FAR = 5%, while KD remains the better choice for FAR = 1% in 5-shot and 10-shot settings. Given that the training uses the train split of the MSWC dataset, it is not surprising to observe that combining KD loss with a task-specific loss can improve the performance when evaluated on the test split of same dataset. However, for the tests on the GSC dataset, KD strategy is consistently the best across the metrics of interest. In this cross-domain setting where the speaker and recording characteristics are different in the test scenario, the exclusion of task-specific losses is critical in preventing overfit on the training conditions. In light of both sets of results, we recommend KD for practical applications requiring low false alarm rates, providing robustness in all settings.

### v2

@@ Abstract
Keyword Spotting plays a critical role in enabling hands-free interaction for battery-powered edge devices. Few-Shot Keyword Spotting (FS-KWS) addresses the scalability and adaptability challenges of traditional systems by enabling recognition of custom keywords with only a few examples. However, existing FS-KWS systems achieve subpar accuracy at desirable false acceptance rates, particularly in resource-constrained edge environments. To address these issues, we propose a training scheme that leverages self-supervised learning models for robust feature extraction, dimensionality reduction, and a teacher-student framework for knowledge distillation. The teacher model, based on Wav2Vec 2.0 is trained using Sub-center ArcFace loss, which enhances inter-class separability and intra-class compactness. To enable efficient deployment on edge devices, we introduce attention-based dimensionality reduction and train a standard lightweight ResNet15 student model. We evaluate the proposed approach on the English portion of the Multilingual Spoken Words Corpus (MSWC) and the Google Speech Commands (GSC) datasets. Notably, the proposed training method improves the 10-shot classification accuracy from 33.6% to 76.9% on 11 classes at 1% false alarm accuracy on the GSC dataset, thus making it significantly better-suited for a real use case scenario.
@@ I Introduction
Keyword Spotting (KWS) enables hands-free interaction with devices by detecting specific spoken commands or wake-words in audio streams. Traditional KWS systems are limited in adaptability and scalability, as they require thousands of training examples for a fixed vocabulary set, and require computation and memory resources beyond the capabilities of embedded edge devices, often battery powered smart sensors. Few-shot keyword spotting systems (FS-KWS) address these challenges by enabling the recognition of custom keywords from only a handful of examples, using deep neural networks up to only a few hundred thousand parameters.
@@
FS-KWS leverages metric learning frameworks allowing the model to readily adapt to new classes (i.e. keywords) by enrolling only a few examples [1]. Notably, the approach of prototypical networks is to train a representation model that processes speech segments into fixed dimensional feature vectors, also referred to as embeddings [2]. The goal is to learn an embedding space in which samples of the same keyword cluster together, while those of different keyword classes are well separated. This is achieved by metric learning objectives such as Triplet loss or Prototypical Loss [3]. For systems that support the open-set setting, users provide a few spoken examples of their keywords during the enrollment phase, from which class prototypes are computed as the mean of the embeddings [4]. The inference decisions are based on the shortest distance between the embedding of the test samples and the enrolled class prototypes. Unfortunately, the leading approaches in this new field achieve low accuracies at desired false acceptance rates [5, 6].
@@
A common approach to boost the performance of supervised speech tasks is to incorporate self-supervised learning (SSL) models, such as Wav2Vec 2.0 [7] or HuBERT [8], which generate rich frame-level features. These SSL models output a feature vector for each frame of speech, resulting in many vectors per the duration of a keyword. Methods such as mean-pooling [9], convolutional encoders [10], or attention-based extraction [11] have been used to obtain lower dimensional embeddings for KWS and other discriminative tasks.
@@
In this paper, we propose a FS-KWS model training scheme as shown in Fig. 1 that combines a pretrained SSL model with an alternative metric learning objective. Specifically, our training scheme leverages knowledge distillation, with the SSL-based teacher network optimized by Sub-center ArcFace (SCAF) loss  [12], as opposed to triplet loss variants. ArcFace  [13], originally developed for face recognition, enforces class separation by adding an angular margin to the softmax classifier. SCAF further enhances this by introducing multiple sub-centers per class and selecting the closest one for each sample, better capturing intra-class variation. These sub-centers allow the model to represent diverse samples within the same class more effectively. ArcFace has recently been applied to audio-related tasks such as emotion recognition [14] and speaker verification [15]. To the best of our knowledge, our work is the first to employ SCAF in the context of audio discrimination for FS-KWS. We further explore training strategies for edge-friendly student models, and benchmark our approach against the leading FS-KWS method on public datasets.
@@
  • Utilization of SCAF for efficient dimensionality reduction of embeddings from a pre-trained SSL model.
@@ II-A The Teacher Representation Model
Figure 2: Proposed Teacher Representation Model with (a) simple pooling encoder, (b) attention encoder dimensionality reduction models
@@
1) Simple Pooling Encoder given in Fig. 2(a) implements a common dimensionality reduction approach, where temporal average pooling is followed by linear projection to obtain lower-dimensional representations.
2) Attention Encoder given in Fig. 2(b) incorporates a Scaled Dot Product Attention [17] mechanism followed by parametric rectified linear unit (PReLU) [18] activation and a one-dimensional convolutional layer (Conv1D). The attention mechanism computes temporal relationships across the input sequence, enabling the model to focus on salient temporal features. The subsequent Conv1D layer summarizes the temporal dimension to a single value through weighted averaging, effectively preserving critical temporal information while reducing computational complexity.
@@
Figure 3: Performance of (a) Teacher models on the MSWC dataset under 1-shot setting for different dimensionality reduction architectures (b) Student models on the MSWC dataset under 1-shot setting (c) Student models on the GSC dataset under 10-shot setting
These dimensionality reduction models are trained using labeled samples with two distinct loss functions: (i) Triplet ($\mathcal{L}_{\text{triplet}}$) and (ii) SCAF ($\mathcal{L}_{\text{subcenter}}$). The experiments carried out to determine the optimal dimensionality reduction model and the loss function are provided in Section III. Owing to space limitations, we refer interested readers to Section II of the supplementary material for definitions of the loss functions.
@@ II-B Edge-friendly Student Representation Model
ResNet-based models are widely utilized as benchmarks in edge-focused keyword spotting (KWS) approaches [19, 20, 21]. Specifically, the ResNet15 model has demonstrated superior performance over edge-friendly candidates in [5]. Our experiments therefore focus on ResNet15, which takes mel-frequency cepstral coefficients (MFCC) as input to further reduce data size, memory and computation, making it ideal for edge hardware. Compared to the teacher representation model, the student ResNet15 has three orders of magnitude fewer parameters (480480k vs. 218218M) and computational requirements (235235M multiply-accumulate operations vs. 63.363.3G).
@@
In this work, we employ mean squared error (MSE) as the distillation loss to align the embeddings of the teacher and student models. While the distillation loss by itself results in decent student models, we further explore optimizing the models via a task-specific loss in Section III, for which we consider the same two alternatives as in training the teacher network: Triplet ($\mathcal{L}_{\text{triplet}}$) and SCAF ($\mathcal{L}_{\text{subcenter}}$).
@@ III-A Datasets
To train the dimensionality reduction network of the teacher, and the edge-friendly student network, we use the train split of the English portion of Multilingual Spoken Words Corpus (MSWC) [23], which contains 5.5 million one-second samples, belonging to 39,000 unique words. We test our approach using the test split of the MSWC dataset, comprised of 700k samples from 8900 unique words, in addition to the Google Speech Commands (GSC) dataset [24], a collection of 100k one-second audio clips belonging to 35 short commands.
@@ III-B Evaluation Protocol
To evaluate the few-shot keyword spotting performance, we adopt the same testing protocol as in [5], both for teacher and student models. During the enrollment phase, $K$ samples for each keyword is fed to the frozen representation model, and the mean of the generated embeddings of each keyword class is computed to generate class prototypes. For inference, the test sample is fed to the feature extractor, and the cosine distance between the test sample embedding and the prototypes of each keyword class is computed. If the closest of these distances is less than a threshold $T$, the test sample is assigned to the class with the nearest prototype, otherwise it is assigned to the “others” class. By changing the threshold $T$, the tradeoff between accuracy levels and false acceptance rate (FAR) of “others” can be analyzed for different $K$-shot scenarios.
@@
For the GSC dataset, we adapt the original KWS-12 test setting [24] for FS-KWS. The task is to classify 11 classes (target keywords: {on, off, left, right, up, down, go, stop, yes, no} and silence), while ideally rejecting the 25 remaining classes as “others”. The silence class includes recordings with various background noises. It should be noted that the inclusion of the silence class makes our evaluation protocol more challenging compared to the baseline work of [5]. For enrollment, $K$ samples per keyword in the training set are used to generate the class prototypes, and the entire test set is evaluated against these. To reduce variance from the few-shot sample selection, each test is repeated 100 times and results are averaged.
@@ III-C Selection of Dimensionality Reduction Model
To identify the optimal teacher model for training the lightweight representation model, we conducted experiments utilizing two dimensionality reduction architectures alongside various loss functions. In our initial experiment, we assessed the performance of the two architectures, both of which were trained using triplet loss. Each model underwent triplet training for 3000 batches, with each batch sampling 512 triplets from randomly selected word samples. The squared normalized Euclidean distance was computed for each pair of samples. We set the triplet margin to 0.5 and the model checkpoint that minimized the loss on the validation set was selected for subsequent testing. As anticipated, the attention-based dimensionality reduction architecture yielded superior results over average pooling. Consequently, we focused our investigation on the selection of loss functions specifically for this architecture. Substituting the loss function with Subcenter ArcFace loss [12] and training the dimensionality reduction architecture for 10 epochs using an angular margin of $m=28.6$, a scale factor of $s=32$, and $K=3$ subcenters per class led to the best results. Repeating training with the final hyperparameters and three different random seeds, validation loss varied by less than 0.5%.
@@
Fig. 3(a) illustrates the performance of three different models on the MSWC dataset in the 1-shot setting. The results indicate that the attention-based dimensionality reduction architecture significantly outperforms others when trained with SCAF loss. Specifically, the average inter-class cosine distance between embedding pairs on the test set increases from 0.89 to 0.93 with the inclusion of the attention encoder, and further to 0.95 when the SCAF loss is applied. Correspondingly, the average intra-class distance decreases from 0.27 to 0.26 and 0.25 respectively. The distributions of these distances are provided in Section III of the supplementary material. For completeness, we trained student models using all three models as teachers. The attention-based architecture with SCAF loss also yielded the best results for the student models. For the scope of this letter, we report results exclusively for student models trained with this teacher model.
@@
TABLE I: Comparison of teacher and student (edge) network performance across MSWC and GSC datasets for {1, 5, 10}-shot settings.
Model Objective # of supports MSWC GSC
$DET_{1\%}$ $DET_{5\%}$ AUROC $ACC_{1\%}$ $ACC_{5\%}$ AUC
Teacher SCAF 1-shot 91.6 97.4 99.3 69.0 77.1 83.4
Edge (Resnet15) Triplet [5] 72.3 88.3 97.5 23.7 52.3 71.1
SCAF 76.8 88.5 92.5 19.9 37.3 56.3
KD 83.7 92.7 98.1 42.9 59.2 73.0
KD + Triplet 86.3 95.4 98.9 29.5 47.6 66.5
KD + SCAF 86.3 93.9 98.3 44.0 59.0 73.7
Teacher SCAF 5-shot 96.8 99.2 99.8 81.6 84.8 89.9
Edge (Resnet15) Triplet [5] 89.8 97.0 99.3 32.6 70.6 84.7
SCAF 89.4 94.6 96.4 40.4 59.0 74.2
KD 95.1 97.9 98.1 69.2 80.5 88.4
KD + Triplet 95.4 98.8 99.6 53.5 71.1 83.6
KD + SCAF 95.1 98.1 99.4 69.5 80.2 88.6
Teacher SCAF 10-shot 96.9 99.3 99.9 82.2 85.2 90.9
Edge (Resnet15) Triplet [5] 91.1 97.6 99.4 33.6 76.4 87.5
SCAF 91.0 95.9 97.1 47.9 65.5 78.1
KD 95.7 98.2 99.5 74.4 83.9 90.2
KD + Triplet 95.5 98.9 99.7 55.8 74.0 86.1
KD + SCAF 96.1 98.4 99.5 76.9 85.2 91.0
III-D Comparative Analysis of Edge Model Training Strategies
@@
In order to systematically evaluate the training strategies, we trained the edge-friendly ResNet15 model using four approaches and compared these to the baseline method [5], wherein the same ResNet15 was trained with Triplet loss (Triplet). We directly used the trained checkpoint from the authors’ repository [25] to evaluate the baseline method with our testing protocol. As in the baseline method, we use the first 10 MFCCs, computed using a frame size of 40 ms, a stride of 20 ms, and applying Hamming windowing for the input to the model. This results in a feature map of size 49x10 for each 1-second test sample.
@@
The first strategy replaces Triplet loss in the baseline with SCAF (SCAF). The second strategy employs knowledge distillation from the selected teacher model discussed in the previous section, with the following configurations: (1) KD: trained solely with $\mathcal{L}_{KD}$ ($\lambda=0$); (2) KD + Triplet: trained with both $\mathcal{L}_{KD}$ and $\mathcal{L}_{\text{triplet}}$ ($\lambda=0.03$), thereby balancing MSE and Triplet losses; (3) KD + SCAF: trained with $\mathcal{L}_{KD}$ and $\mathcal{L}_{\text{subcenter}}$ ($\lambda=0.0003$). The selected $\lambda$ values are determined empirically. Each student model was trained three times with different seeds; validation losses and AUC values in Table I varied by less than 0.6% in all cases.
@@
Fig. 3(c) shows accuracy versus FAR for the GSC dataset for the 10-shot setting. The KD model consistently outperforms all other configurations at FAR = 1% and 5%. The baseline Triplet model ranks second at FAR = 5% but drops sharply at lower FARs. This is primarily due to the inclusion of the silence class, as we observed in keyword-specific analysis not shown in this letter. We hypothesize that knowledge distillation from the pre-trained SSL model is key to robustness under unseen conditions.
A comprehensive evaluation of the training strategies is shown in Table I, which includes the best performing teacher model and all edge-friendly model training alternatives. On the MSWC dataset, KD + Triplet and KD + SCAF strategies show marginally better performance than KD alone with higher AUC and detection rates at the given FARs. Since the training uses the train split of the MSWC dataset, it is not surprising to observe that combining KD loss with a task-specific loss can improve the performance when evaluated on the test split of same dataset. However, for the tests on the GSC dataset, the results are mixed. While the KD + Triplet strategy performs significantly worse than KD, KD + SCAF is consistently better than KD for the $ACC_{1\%}$ and $AUC$ metrics. In light of these results, for this cross-domain setting where the speaker and recording characteristics are different in the test scenario, the inclusion of task-specific losses during training should be carefully considered. While the triplet loss evidently introduces overfit on the training conditions, incorporating the SCAF loss maintains the cross-domain performance. Therefore, we recommend KD + SCAF for practical applications requiring low false alarm rates, providing robustness in all settings.